OpenStudy (anonymous):

Solve for y: y^2 + 12y + 11 = 0. A. y = 1 or -11 B. y = -1 or -11 C. y = -1 or 11 D. None of the above

OpenStudy (dls):

\[\LARGE y^2+11y+1+11=0\] \[\LARGE y(y+11)+1(y+11)=0\] \[\LARGE (y+1)(y+11)=0\] This means that..?

OpenStudy (anonymous):

It is none of the above?

OpenStudy (anonymous):

B

OpenStudy (dls):

\[\LARGE y^2+11y+1y+11=0\] \[\LARGE y(y+11)+1(y+11)=0\] \[\LARGE (y+1)(y+11)=0\] sorry typing error in 1st line* and No. Think again?

OpenStudy (anonymous):

Y=-1 x=-11

OpenStudy (dls):

Yup.

OpenStudy (dls):

where did x come from o_O